Here is another one,
solve each system by the addition method.
2x=2-y
3x+y=-1
Help,I have no clue.

Answer by suresh(20)

The given equations are,
2x = 2 - y------------>(1)
3x + y = -1----------->(2)
Eqn (1) becomes, -2x - y = -2 ------->(3)
Adding (1) and (3),
x = - 3.
Plug x = -3 in (1), we have 2 ( -3 ) = 2 - y
-6 = 2 - y
-y = - 8.
y = 8
Hence x = -3 and y = 8